The expression(-1+i)^3 is equivalent to
(1) -3i
(2) -2-2i
(3) -1-i
(4) 2+2i
----------------------------------------------------------------------------------
please choose the correct answer choice and explain how to get the answer
Thank you

Answer by Alan3354(21550)

The expression(-1+i)^3 is equivalent to
(1) -3i
(2) -2-2i
(3) -1-i
(4) 2+2i
-----------
(-1+i)^2 = 1-1 - 2i = -2i
(-1+i)^3 = -2i*(-1+i) = 2i + 2
= 2+2i --> #4
----------------
-1+i = sqrt(2)cis135
(-1+i)^3 = 2sqrt(2)cis405
= 2sqrt(2)cis45
= 2sqrt(2)*(cos(45) + isin(45))
= 2 + 2i Same answer